# Benchmarking Optimizers for MLPs in Tabular Deep Learning

### 16 Apr 2026 v1

@@ 1 Introduction
    We find the Muon optimizer (Jordan et al., 2024b) to consistently outperform AdamW for both plain MLPs and modern MLP-based architectures. The better performance of Muon comes at the cost of slower training, with an average slowdown of $3\times$ for hyperparameter tuning across all datasets.
@@ 3 Experiments
In this section, we evaluate how optimizer choice affects the generalization performance of MLP-based models in supervised tabular learning under a unified experimental protocol. We structure our experiments as follows. First, we provide a complete overview of our tuning and evaluation setup (§3.1). We then present an extensive empirical evaluation of a wide range of optimizers and optimization techniques for the MLP baseline (§3.2). In the last subsection, we evaluate the most performant and effective optimizer variants for a broader set of MLP-based state-of-the-art models (§3.3).
@@ 4 Limitations
Our study focuses specifically on MLP-based architectures for supervised learning on tabular data. We do not cover tabular foundation models, retrieval-based methods, or other non-MLP paradigms, where optimizer behavior may differ. Our results are also purely empirical: understanding why methods such as Muon help in this setting is an important direction for future work. Finally, Muon is measurably slower than AdamW in this setting. Across the 17 MLP tuning runs, Muon required about $3.03\times$ more tuning time than AdamW on average (46.9h vs 15.5h total for all 17 datasets), while AdamW with EMA required about $1.29\times$ more (20.0h total). This extra cost may matter in more compute-constrained use cases.

### 17 Apr 2026 v2

@@ 1 Introduction
    We find the Muon optimizer (Jordan et al., 2024b) to consistently outperform AdamW for both plain MLPs and modern MLP-based architectures. The better performance of Muon comes at the cost of slower training; for stronger models, the efficiency overhead is less pronounced.
@@ 3 Experiments
In this section, we evaluate how optimizer choice affects the generalization performance of MLP-based models in supervised tabular learning under a unified experimental protocol. We structure our experiments as follows. First, we provide a complete overview of our tuning and evaluation setup (§3.1). We then present an extensive empirical evaluation of a wide range of optimizers and optimization techniques for the MLP baseline (§3.2). Then, we evaluate the most performant and effective optimizer variants for a broader set of MLP-based state-of-the-art models (§3.3). Finally, we quantify the training efficiency overhead coming from the best optimizers (§3.4).
@@ 3.3 Muon and EMA for SoTA MLP-based Models
3.4 Training Efficiency Overhead
In Table 2, we quantify the training efficiency overhead coming from using Muon and AdamW with EMA instead of the vanilla AdamW. Notably, the picture is different for simpler MLPs and for the more powerful TabM. First, for TabM, the relative efficiency overhead associated with Muon is generally less prounounced. Second, for MLP, AdamW with EMA is noticeably more efficient than Muon, which is not the case for TabM.
Table 2: Task performance and training efficiency overhead across multiple model families and optimizers. EMA denotes AdamW with EMA. The top row reports the mean improvement relative to MLP trained with AdamW ($\Delta_{\mathrm{Score}}$ is defined in subsection 3.1). The bottom row reports the hyperparameter tuning time overhead relative to AdamW. For both metrics, their mean values across all datasets are reported.
MLP MLP TabM $_{\text{Packed}} TabM$
Metric EMA Muon EMA Muon EMA Muon EMA Muon EMA Muon
$\Delta_{\mathrm{Score}}$ over MLP[AdamW] 0.66 0.32 1.42 1.56 1.75 1.71 1.31 1.45 2.21 2.57
[…]
@@ 4 Limitations
Our study focuses specifically on MLP-based architectures for supervised learning on tabular data. We do not cover tabular foundation models, retrieval-based methods, or other non-MLP paradigms, where optimizer behavior may differ. Our results are also purely empirical: understanding why methods such as Muon help in this setting is an important direction for future work. Finally, replacing AdamW with Muon results in slower training, though the overhead for more powerful models is less noticeable.